OpenStudy (nottim):

Solve the following equation, xe[2,0pi]. tan^3x-9tan^2x+17tanx-6=0

OpenStudy (nottim):

Well I tried putting x=1, I got an error.

OpenStudy (anonymous):

\[x \approx3.14159 n+1.10715\] \[x \approx3.14159 n+0.429998\] \[x \approx3.14159 n+1.4191\]

myininaya (myininaya):

first see if you can solve this \[u^3-9u^2+17u-6=0\]

OpenStudy (jamesj):

Thanks myin; that's the way.

myininaya (myininaya):

if you try u=2, you will see you have one solution

OpenStudy (anonymous):

thats what i wasn gonna say replace tan with a variable

myininaya (myininaya):

then using synthetic division you can find the other 2 solutions

OpenStudy (nottim):

I dunno what that is.

myininaya (myininaya):

synethic division+quadratic formula (or perhaps factoring if we have other rational zeros)

OpenStudy (nubeer):

u can also urs long division to get factors

OpenStudy (nottim):

Wait a sec guys, I'm trying to understand this all...

OpenStudy (jamesj):

A fancy way of saying that as u = 2 is a solution, (u-2) must be a factor of the polynomial. Hence you can write \[ u^3−9u^2+17^u−6 = (u-2)q(u) \] for some second order polynomial

OpenStudy (jamesj):

make that 17u, not 17^u

myininaya (myininaya):

lol

OpenStudy (nottim):

What happened to tan?

OpenStudy (jamesj):

i'll let the other chefs take over

myininaya (myininaya):

we will get back to the tan business after we solve that polynomial=0 for u

OpenStudy (nottim):

Ok, so I'm doing the whole F(#) thing at this point right?

myininaya (myininaya):

what?

myininaya (myininaya):

what does F(#) mean?

OpenStudy (nottim):

Looking for the value of x that makes everything 0.

OpenStudy (anonymous):

u = 2 u = 1/2 (7-sqrt(37)) u = 1/2 (7+sqrt(37))

OpenStudy (anonymous):

If you think how myininaya magically found that \( u=2 \) there is a analytical way of doing it which is known as Rational root test.

myininaya (myininaya):

right fool

OpenStudy (nottim):

What if I got an error on my calculator?

OpenStudy (nottim):

Useless its tanx (as an entirety) that I using....

OpenStudy (anonymous):

Your trying to plug number right into the equation?

myininaya (myininaya):

i went through the possible rational zeros and found a rational zero i didn't go through all possible rational zeros as i only needed one so i can just mess with a quadratic

OpenStudy (nottim):

Lagrange, yes...

OpenStudy (nottim):

Not the right idea huh?

OpenStudy (anonymous):

You need to NotThink a bit.

OpenStudy (nottim):

...ok? I do that a lot.

OpenStudy (anonymous):

Follow those steps that myin showed, youll end with the right answers

OpenStudy (nottim):

I'm just looking at this and my notes, and I have not a clue.

OpenStudy (nottim):

Is u=tanx?

myininaya (myininaya):

yes

OpenStudy (nottim):

Ok.

OpenStudy (anonymous):

it's more like Let u = tan x then form a polynomial in u, factor it and then substitute back the u =tan x and then solve it ..

OpenStudy (nottim):

ERkk.

OpenStudy (nottim):

Imma do this slowly... If yo ugotta do something while you help me, do it now as I work this thru...

OpenStudy (nottim):

IS factor Thereom involved?

OpenStudy (paxpolaris):

yes...if u-2 is a factor of P(u) ...P(2)=0

OpenStudy (paxpolaris):

Let tan (x) = u \[u ^{3}-9u ^{2}+17u-6=0\] \[\left( u-2 \right)\left( u^2 - 7u -3\right) =0\]

OpenStudy (nottim):

wait wait waitt.... And that's also not how I was taught (or shown in the notes). I'll try it out after I'm done my thing tho/.

OpenStudy (nottim):

Sorry about that. My internet went wonky.

OpenStudy (nottim):

And now I divide the thing by u-2 right?

myininaya (myininaya):

u^2-7u+3 ____________________ u-2| u^3-9u^2+17u-6 -(u^3-2u^2) ------------------- -7u^2+17u-6 -(-7u^2+14u) --------------- 3u-6 -(3u-6) -------- 0

OpenStudy (nottim):

ok

myininaya (myininaya):

pax did i make a mistake?

myininaya (myininaya):

our factorizations are a lttle different i have \[\left( u-2 \right)\left( u^2 - 7u +3\right) =0\]

OpenStudy (anonymous):

myininaya, I wish If I can give you more medals for your patience.

OpenStudy (nottim):

Lol.

myininaya (myininaya):

are you talking about me and my division above? lol

OpenStudy (nottim):

I think just waiting for me to finish up my stuff here. Thanks by the way.

myininaya (myininaya):

oh i had a thing i was doing lol

OpenStudy (paxpolaris):

sorry u r right. it's +3

OpenStudy (nottim):

ITs +3?

myininaya (myininaya):

just making sure when it gets late i tend to screw up on little things

myininaya (myininaya):

like sometimes i don't know how to add or multiply sometimes

myininaya (myininaya):

we were talking about this: \[\left( u-2 \right)\left( u^2 - 7u +3\right) =0 \]

OpenStudy (nottim):

ok. that scared me for a sec.

OpenStudy (anonymous):

yep myininaya I was talking about that :D

myininaya (myininaya):

lol

myininaya (myininaya):

so we know we have u=2 we also have \[u=\frac{7 \pm \sqrt{49-4(3)}}{2}\] don't forget to replace u with tan(x) and solve for x

OpenStudy (nottim):

Finally, I'm done the division.

myininaya (myininaya):

lol

myininaya (myininaya):

good

myininaya (myininaya):

you have \[u-2=0 \text{ or } u^2-7u+3=0\]

myininaya (myininaya):

we need to use the quadratic formula to solve the second equation

OpenStudy (nottim):

I hope I remember this for the test.

myininaya (myininaya):

so this is trig class?

myininaya (myininaya):

i could be wrong but this is too long to put on a test

myininaya (myininaya):

there might be a similar problem he/she puts there

myininaya (myininaya):

but shorter

OpenStudy (nottim):

Whoops. I didn't type that...Double post or something.

myininaya (myininaya):

i could be wrong

OpenStudy (anonymous):

sometimes teachers get problems from special books

myininaya (myininaya):

I'm sleepy guys I'm sure if you need more help NotTim (who we decided was not myininaya earlier ;)lol) , fool and pax will continue to help unless they are are going.

OpenStudy (nottim):

K. Thanks man.

myininaya (myininaya):

Do I look like a man to you?

myininaya (myininaya):

Please say no

myininaya (myininaya):

lol

OpenStudy (nottim):

Are you?

OpenStudy (nottim):

I actually have bad vision and human identification problems.

OpenStudy (nottim):

When I'm walking around I will misidentify everyone as the same person.

myininaya (myininaya):

Ok. Its cool. No worries.

OpenStudy (nottim):

any1 know what I should do at (m-2)(m^2-7m+3)?

OpenStudy (nottim):

For m^2-7m+3, do I have to do the quadratic formula?

OpenStudy (nottim):

I guess I already know what to do with m-2, so I'm fine there.

OpenStudy (nottim):

ERm, are you people just gunna lurk there? It's fine, just....

OpenStudy (anonymous):

find all places that u can be 0, then, recalling that u = tanx, solve for x.

OpenStudy (anonymous):

you already found one solution is u=2, you can find the others by using the quadratic formula on the quadratic part of your factoring.

OpenStudy (nottim):

Ok.

OpenStudy (nottim):

Hey, if one of my answers is 6.54, and I wanna get the other angle with it, what do i Do?

OpenStudy (nottim):

Hi?

OpenStudy (nottim):

Um, sorry if I sound abrupt, but could anyone please answer my most recent question?